# MBSS-T1: Model-Based Self-Supervised Motion Correction for Robust Cardiac T1 Mapping

### v2

@@ 3.1 Model Architecture
Figure 2 presents our overall architecture. Our network comprises two U-Net-inspired encoder-decoder modules that operate simultaneously. The network has three main components:
Motion Estimation Component: The upper module, an extension of the VoxelMorph architecture, functions as a multi-image deformable image registration module. It dynamically adjusts the number of registration heads based on the number of recovery inversion time points (). It outputs a specific deformation field and registered image for each timestamp . This module leverages the deformation field prediction layers from [2], and the integration layer from [3].
Parametric Mapping Component: The bottom module predicts the parameters of the qMRI signal relaxation model (T1 and ) and forwards these parameters to the signal generation layer. This layer predicts a sequence of synthetic images following the Inversion Recovery T1 Relaxation Equation (Eq. 1). These images are anatomically aligned according to the correct signal decay model, as each image prediction utilizes the same T1 and parameters, with the only difference being the timestamp of each image.
Both encoder-decoder modules follow a U-Net-like structure, including convolutional and downsampling layers and skip connections that establish links between each module’s encoder and decoder.
@@
Segmentation and Uncertainty Component: To ensure the anatomical correctness of the aligned images, we added anatomical constraints to the registration process. First, nnU-Net [17] (other segmentation networks can also be used) is utilized as a pretrained network to predict the myocardium in every unregistered frame. Second, an uncertainty block is employed to retain only the segmentations with high confidence in their correctness. This block receives the masks and softmax maps generated by the nnU-Net output. It retains only those masks that meet two criteria: (1) a required percentage of the myocardial segmentation pixels must have a score greater than a confidence threshold , and (2) the segmentation mask must consist of exactly one connected component. Here, denotes the softmax probability at pixel in the -th segmentation, and is the corresponding segmentation mask. Let be the set of indices of segmentations that satisfy both of these criteria:
Next, the masks are aligned using the deformation fields from the Motion Estimation Component to obtain deformed segmentations. During registration, the Dice score is maximized between the deformed myocardium segmentations to ensure that the myocardiums are properly registered. We define a reference image index as the time point corresponding to the segmentation with the highest confidence score:
@@
This equation selects the index where the average confidence across the segmentation is the highest among the segmentations in the set .
@@ 3.2 Loss Functions
The first term () addresses the differences between the images predicted by the model’s decoder and the deformed images. To achieve this, we compute the mean squared error (MSE) between the registered images and the synthetic images :
Here, denotes the images generated using the signal model equation (Eq. 1), while the registered images are obtained from the Parametric Mapping component. This term enforces the creation of deformation fields that are physically plausible by ensuring consistency with the T1 signal relaxation model.
@@ 4.1.1 Free Breathing - STONE Dataset
A publicly available myocardial T1 mapping dataset [4, 5] consists of 210 subjects, 134 males and 76 females, aged years. All subjects were diagnosed with or suspected of having cardiovascular diseases. The imaging was performed on a 1.5T MRI scanner (Philips Achieva) equipped with a 32-channel cardiac coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping sequence (STONE) [30]. The acquisition parameters were as follows: field of view (FOV) [mm2] and voxel size [mm3]. Five slices were captured from the base to the apex in the short-axis view for each subject at 11 distinct time points. Additionally, the dataset included manual expert segmentations of the myocardium [4]. The images were resized to pixels per time point, and min-max normalization was applied to the entire sequence to standardize the image intensities.
@@ 4.1.2 Free Breathing and Breath Hold - MOLLI Dataset
Data was acquired on a 3T Siemens MRI system (PRISMA, Siemens Healthineers, Erlangen, Germany) equipped with a 32-channel body coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping MOLLI sequence (MyoMaps, SIEMENS Healthcare). The acquisition parameters were as follows: field of view (FOV) [mm2], flip angle of 35°, and an 8 mm gap between slices. Five slices were captured from the base to the apex in the short-axis view for each subject at eight distinct time points. The dataset consists of cardiac images obtained from 19 patients undergoing two scans: one during breath-hold and another during free breathing. Additionally, extra breath-hold and free-breathing scans were performed for five patients for test-retest analysis.
@@ 4.2 Experiments
Experimental setup: Our MBSS-T1 method was compared against two variations: a multi-image registration model using a mutual-information-based loss function (REG-MI) [12], and PCMC-T1, which incorporates physical constraints into the motion correction process within an unsupervised framework [13]. Additionally, two state-of-the-art deep learning algorithms for medical image registration were included in the comparison: pairwise probabilistic diffeomorphic VoxelMorph with a mutual-information-based loss [3] and pairwise SynthMorph [15]. For the self-supervised approaches, MBSS-T1 and SynthMorph, the entire STONE dataset was utilized for inference since no training phase was required. For the unsupervised approaches (PCMC-T1, REG-MI, VoxelMorph) and the nnUNet pretrained segmentation network (used for MBSS-T1), a 5-fold cross-validation strategy was implemented. In each fold, the 210 subjects were divided into 80% for training and 20% for testing. T1 maps generated directly from the acquired images were also used for comparison by applying a standard T1 mapping procedure. A T1 relaxation curve was fitted to the MRI data using the conventional inversion recovery model without any pre-motion correction techniques. Fitting was performed per pixel across the series of T1-weighted images, resulting in maps representing the T1 relaxation times. These maps, referred to as ’w.o motion correction,’ serve as a baseline to evaluate the impact of motion correction on the accuracy and quality of the T1 mapping. The T1 maps generated by our MBSS-T1 model were quantitatively evaluated against those produced both before (’w.o motion correction’) and after deep-learning-based image registration as a pre-processing step. The evaluation metrics used were the of the model fit to the observed data in the myocardium, the Dice score, and the Hausdorff distance values of the myocardium segmentations.
@@
Experimental Setup: The objective of this experiment was to evaluate the effectiveness of the MBSS-T1 method in accounting for motion during free-breathing and breath-hold MRI scans using the MOLLI sequence. We utilized the pre-trained nnUNet model trained on the STONE dataset, as described in the previous section. To make the MBSS-T1 model compatible with the MOLLI sequence, we modified the MBSS-T1 model by adjusting the number of heads in the network output from 11 to 8. The performance of three motion-correction methods was compared: (1) without motion correction, (2) Siemens MyoMaps, and (3) MBSS-T1. For Siemens MyoMaps, we used the motion-corrected T1 maps produced by the scanner’s image processing software.
@@
Clinical Impact: As in Exp 1, the clinical impact of our method was assessed through a semi-quantitative ranking of the T1 maps for the presence of motion artifacts. This assessment was conducted on free-breathing and breath-hold scans to determine how effectively each method mitigated motion artifacts. The primary evaluation metric was a clinical score assigned by an experienced cardiac radiologist with five years of experience based on the clarity and quality of the T1 maps produced by each method.
@@
4.2.3 Exp. 3: Consistency of T1 Mapping with Motion Correction in Test/ReTest MOLLI Dataset
@@
Experimental Setup and Objective: This experiment aimed to assess the consistency and reproducibility of T1 mapping when motion correction is applied. This analysis aimed to evaluate whether motion correction techniques improve the reliability of myocardial T1 measurements across repeated scans, particularly in the presence of motion. The subset included data from five patients undergoing two free-breathing and two breath-hold scans. The goal was to compare the consistency of T1 measurements across these scans using different motion correction methods.
@@
Quantitative Evaluation: To quantitatively assess the reproducibility of T1 measurements, we calculated the intraclass correlation coefficient (ICC3) [20] for mean myocardial T1 values across three different methods: without motion correction, Siemens MyoMaps, and MBSS-T1. The analysis was conducted using data from five patients in the MOLLI dataset, each of whom underwent two free-breathing and two breath-hold scans. For each scan and method, mean myocardial T1 values were computed for four regions of interest (ROIs) – upper, right, lower, and left – on the mid-slice of the T1 map. The ICC3 metric was then calculated for all paired test-retest ROIs to assess the agreement between repeated measurements. A high ICC3 value indicates strong agreement between the test and retest measurements, suggesting robustness and reproducibility in quantifying myocardial T1 values. This metric provides valuable insights into the reliability and consistency of T1 measurements across different methods and scanning conditions. A comprehensive understanding of each method’s performance in clinical practice can be achieved by quantifying the agreement between test and retest measurements across different ROIs and scanning conditions. This analysis also enables the evaluation of the effectiveness of motion correction techniques in improving measurement accuracy and reproducibility.
@@ 5 Results
Clinical impact: Figure 3 presents several representative cases. The maps generated by the MBSS-T1 method demonstrate the highest quality, as evidenced by superior metrics. The rightmost column of Table 1 summarizes the clinical impact assessment results for the MBSS-T1 method, which achieved the highest quality score of . This score reflects an improvement of approximately 10% over the following best method, PCMC-T1, which had a score of . The difference in radiologist grading was statistically significant (p ). Furthermore, as demonstrated in Fig. 5, the MBSS-T1 method achieves a closer fit to the expected magnetization curve (with ) and reduces artifacts, which may contribute to improved diagnostic confidence. These results suggest the MBSS-T1 method’s potential utility in clinical settings where motion artifacts can impact the accuracy of T1 mapping.
The results of Exp 2 indicate that the MBSS-T1 method improves the quality of T1 maps by effectively reducing motion artifacts, as shown by the highest clinical scores in both breath-hold (BH) and free-breathing (FB) scans. Specifically, MBSS-T1 achieved mean scores of for BH and for FB scans (Table 2). These scores represent an improvement of approximately 5% in BH and 6% in FB scans compared to Siemens Myomap, which had scores of and , respectively.
The images in Fig. 4 illustrate the reduction in motion-related distortions achieved by the MBSS-T1 method. This improved motion correction results in more precise and reliable T1 maps with fewer artifacts.
5.3 Exp. 3: Consistency of T1 Mapping with Motion Correction in Test/ReTest MOLLI Dataset
The results in Table 3 show the ICC3 metrics for mean and median myocardial T1 values across three different motion correction methods: without motion correction, Siemens Myomap, and MBSS-T1.
For both mean and median myocardial T1 values, the MBSS-T1 method consistently exhibited higher ICC3 values than the other methods. Specifically, for the mean T1 values, the MBSS-T1 method attained ICC3 values of for free-breathing (FB), for breath-hold (BH), and for combined FB and BH scans. In contrast, the w.o motion correction and Siemens Myomap methods displayed significantly lower ICC3 values, suggesting less agreement between test and retest measurements. Similarly, for the median T1 values, the MBSS-T1 method achieved ICC3 values of (FB), (BH), and (FB+BH), surpassing the other methods.
These findings suggest that the MBSS-T1 method offers more reliable and consistent myocardial T1 measurements than the other methods. However, it is important to note that while the MBSS-T1 method shows the highest ICC3 values, the scores indicate only moderate reliability. Although the MBSS-T1 method improves the robustness and reproducibility of T1 mapping under free-breathing and breath-hold conditions, further improvements are needed to achieve high reliability.
@@
| Method | Clinical Score | |

### v3

@@ 3.1 Model Architecture
Figure 2 presents our overall architecture. Our system comprises a pre-trained nnUNet for myocardium segmentation and two U-Net-inspired encoder-decoder modules that operate simultaneously to predict the deformation fields and the signal recovery model parameters. We describe each component in detail below.
@@
Segmentation and Confidence Component: First, an nnUNet segmentation network [19] pre-trained on the STONE dataset [4, 5] is utilized to predict the myocardium segmentation for each frame. This network was trained using a 5-fold cross-validation setup to avoid data leakage. The output of the nnUNet includes both the segmentation of the myocardium and a pixelwise confidence score for the segmentation. Next, a confidence block is employed to retain only the segmentations with high confidence scores. This block receives the masks and confidence maps generated by the nnU-Net output. It retains only those masks that meet two criteria: (1) a required percentage of the myocardial segmentation pixels must have a score greater than a confidence threshold , and (2) the segmentation mask must consist of exactly one connected component. Here, denotes the softmax probability at pixel in the -th segmentation, and is the corresponding segmentation mask. Let be the set of indices of segmentations that satisfy both of these criteria:
Next, we define a reference image index as the time point corresponding to the segmentation with the highest confidence score:
@@
This equation selects the index where the average confidence across the segmentation is the highest among the segmentations in the set . This reference image will serve as the fixed image for the registration network. In addition, we define the extended myocardium mask as the bitwise OR of the segmentations in the set .
Motion Estimation Component: The upper module, an extension of the VoxelMorph architecture, functions as a multi-image deformable image registration module. It dynamically adjusts the number of registration heads based on the number of the magnitude images () reconstructed from the acquired data. Its input consists of a concatenation of the reference image , selected as the frame with the highest confidence score from the segmentation and confidence component output, the remaining images as the moving images, and the segmentations retained in the set , which includes only those with high confidence. The output of this module is a set of deformation fields between the moving images and the reference image as well as the deformed images for each time point . This module leverages the deformation field prediction layers from [2], and the integration layer from [3] to encourage diffeomorphism.
Next, the segmentation masks associated with each time point are aligned using the corresponding deformation fields to obtain deformed segmentations.
Parametric Mapping Component: The bottom module takes the same input as the Motion Estimation Component and predicts the parameters of the signal recovery model, forwarding these parameters to the signal generation layer. This layer predicts a sequence of synthetic images following the signal recovery equation suitable for the sequence used for the acquisition (i.e. Eq. 1 for the MOLLI sequence and Eq. 3 for the STONE sequence). These images are anatomically aligned according to the correct signal recovery model, as each image prediction utilizes the same signal recovery model parameters, with the only difference being the times point of each image.
@@ 3.2 Loss Functions
The first term () addresses the differences between the images synthesized by the signal recovery layer equation (i.e. Eq. 1 for the MOLLI sequence and Eq. 3 for the STONE sequence) from the parameters predicted by the parametric mapping component and the deformed images. To achieve this, we compute the mean squared error (MSE) between the registered images and the synthetic images after applying the mask :
Here, denotes the images generated using the signal recovery equation (Eq. 1 or Eq. 3) using the predicted model parameters (), while the registered images are obtained by applying the deformation fields predicted by the motion estimation component to the acquired magnitude images . This term enforces the creation of deformation fields that are physically plausible by ensuring consistency with the T1 signal recovery model.
@@ 3.3 Implementation details
Our network architecture includes two UNet backbones. Each backbone employs 2D convolutions in both the encoder and decoder stages, using a kernel size of 3, stride of 1, and padding of 1. The encoder consists of three convolutional layers with [16, 32, 32] filters, each followed by a LeakyReLU activation function with a negative slope of 0.2. To progressively reduce spatial dimensions, a max-pooling operation with a kernel size of 2, stride of 2, padding of 0, and dilation of 1 is applied after each convolutional layer.
The decoder contains four convolutional layers with 32 filters, each followed by a LeakyReLU activation with a slope of 0.2, along with an upsampling operation with a scaling factor of 2. Subsequently, three additional convolutional layers with [32, 16, 16] filters are applied, each followed by a LeakyReLU activation function.
In the registration heads, each head comprises two convolutional layers with [4, 4] filters, each followed by a LeakyReLU activation with a slope of 0.2. This is followed by another convolutional layer with 2 filters, and a resize block with a scaling factor of 0.5. The velocity field is predicted at every two pixels and processed through seven integration steps, followed by a resizing operation with a factor of 2.
In the parameter heads, each head consists of two convolutional layers with a single filter, each followed by a LeakyReLU activation with a slope of 0.2. This setup ensures the integration of both registration and parameter estimation components into the network’s architecture.
@@ 4.1.1 Free Breathing - STONE Dataset
A publicly available myocardial T1 mapping dataset [4, 5] consists of 210 subjects, 134 males and 76 females, aged years. All subjects were diagnosed with or suspected of having cardiovascular diseases. The imaging was performed on a 1.5T MRI scanner (Philips Achieva) equipped with a 32-channel cardiac coil, using an ECG-triggered, free-breathing, respiratory-navigated, slice-interleaved cardiac T1 mapping sequence (STONE) [37]. The acquisition parameters were as follows: field of view (FOV) [mm2] and voxel size [mm3]. Five slices were captured from the base to the apex in the short-axis view for each subject at 11 distinct time points. Additionally, the dataset included manual expert segmentations of the myocardium [4]. The images were resized to pixels per time point, and min-max normalization was applied to the entire sequence to standardize the image intensities.
@@ 4.1.2 Free Breathing and Breath Hold - MOLLI Dataset
Data was acquired on a 3T Siemens MRI system (PRISMA, Siemens Healthineers, Erlangen, Germany) equipped with a 32-channel body coil, using an ECG-triggered, slice-interleaved cardiac T1 mapping MOLLI sequence (MyoMaps, SIEMENS Healthcare). The acquisition parameters were as follows: field of view (FOV) [mm2], flip angle of 35∘, and slice thickness of 8 mm. Five slices were captured from the base to the apex in the short-axis view for each subject at eight distinct time points. The dataset consists of cardiac images obtained from 19 patients undergoing two scans: one during breath-hold and another during free breathing. The data was acquired without any respiratory navigation. Additionally, extra breath-hold and free-breathing scans were performed for five patients for test-retest analysis. Pre-processing was performed following the same procedure described above for the STONE dataset.
@@ 4.2 Experiments
Experimental setup: We used an MBSS-T1 version (MBSS-T1) with a suitable model parameters estimation module (i.e Eq. 3). We compared our MBSS-T1 against two variations: a multi-image registration model using a mutual-information-based loss function (REG-MI) [14], and PCMC-T1, which incorporates physical constraints into the motion correction process within an unsupervised framework [15]. Additionally, two state-of-the-art deep learning algorithms for medical image registration were included in the comparison: pairwise probabilistic diffeomorphic VoxelMorph with a mutual-information-based loss [3] and pairwise SynthMorph [17].
For the unsupervised approaches (PCMC-T1, REG-MI, VoxelMorph) and the nnUNet pretrained segmentation network (used for MBSS-T1), a 5-fold cross-validation strategy was implemented. In each fold, the 210 subjects were divided into 80% for training and 20% for testing. For each fold, the VoxelMorph model was trained in an unsupervised manner, with the fixed and moving images randomly selected from the training data. The reported results represent the aggregation of the test data across the entire dataset.
The SynthMorph training procedure does not require any training data and was therefore applied directly to the entire STONE dataset. Similarly, MBSS-T1 was applied directly to the entire STONE dataset without any prior training.
The reference time point for registration was automatically determined by the MBSS-T1 approach and was subsequently set to the first time point for all other methods.
T1 maps generated directly from the acquired images were also used for comparison by applying a standard cardiac T1 mapping procedure. A T1 relaxation curve was fitted to the MRI data using the conventional inversion recovery model without any pre-motion correction techniques. Fitting was performed per pixel across the series of T1-weighted images, resulting in maps representing the T1 relaxation times. These maps, referred to as ’w.o motion correction,’ serve as a baseline to evaluate the impact of motion correction on the accuracy and quality of the cardiac T1 mapping.
The T1 maps generated by our MBSS-T1 model were quantitatively evaluated against those produced both before (‘w.o motion correction’) and after deep-learning-based image registration as a pre-processing step. The evaluation metrics used were the of the model fit to the observed data in the myocardium, the Dice score, and the Hausdorff distance values of the myocardium segmentations. In addition, we evaluated the capacity of the different methods to generate diffeomorphic deformation fields by assessing the determinants of the deformation fields Jacobians for each method.
@@
Experimental Setup: The objective of this experiment was to evaluate the effectiveness of the MBSS-T1 method in accounting for motion during free-breathing and breath-hold MRI scans using the MOLLI sequence. We utilized the pre-trained nnUNet model trained on the STONE dataset, as described in the previous section. To make the MBSS-T1 model compatible with the MOLLI sequence (MBSS-T1), we modified the MBSS-T1 model by adjusting the number of heads in the network output from 11 to 8 and replacing the model fitting layers with layers that match the MOLLI sequence signal recovery model and T1 correction equation(Eq. 1, 2). The performance of four motion-correction methods was compared: (1) without motion correction, (2) Siemens MyoMaps with motion correction [38], (3) Synthmorph [17] and (4) MBSS-T1.
The reference time point for registration was automatically determined by the MBSS-T1 approach and was subsequently set to the first time point for Synthmorph.
For Siemens MyoMaps, we used the motion-corrected T1 maps produced by the scanner’s image processing software. We quantitatively evaluate our method against comparative approaches using three key metrics: the coefficient of determination () for model fit to the observed myocardial data, the Dice score, and the Hausdorff distance for myocardial segmentations. Since manual myocardial annotations are unavailable for our MOLLI dataset, we applied a pretrained segmentation network to each frame individually and computed the segmentation metrics both before and after motion correction. These metrics were not measured for Siemens MyoMap, as we do not have access to the registered T1-weighted images.
@@
Clinical Impact: As in Exp. 1, the clinical impact of our method was assessed through a semi-quantitative ranking of the T1 maps for the presence of motion artifacts. This assessment was conducted on free-breathing and breath-hold scans to determine how effectively each method mitigated motion artifacts. The primary evaluation metric was a clinical score assigned by an experienced cardiac radiologist with five years of experience based on the clarity and quality of the T1 maps produced by each method.
@@
4.2.3 Exp. 3: Consistency of cardiac T1 Mapping with Motion Correction in Test/ReTest MOLLI Dataset
@@
Experimental Setup and Objective: This experiment aimed to assess the consistency and reproducibility of cardiac T1 mapping when motion correction is applied. This analysis aimed to evaluate whether motion correction techniques improve the reliability of myocardial T1 measurements across repeated scans, particularly in the presence of motion. The subset included data from five patients undergoing two free-breathing and two breath-hold scans. It is important to note that the test and retest images are not aligned, as the subject left the room for a 30-minute break between scans. As a result, the images are not directly comparable on a pixel level. Instead, comparisons between the test and retest data are conducted at a higher level of abstraction, focusing on regional metrics rather than pixel-level correspondence.
@@
Quantitative Evaluation: To quantitatively assess the reproducibility of T1 measurements, we calculated the intraclass correlation coefficient (ICC3) [23] for mean myocardial T1 values across three different methods: without motion correction, Siemens MyoMaps, and MBSS-T1. The analysis was conducted using data from five patients in the MOLLI dataset, each of whom underwent two free-breathing and two breath-hold scans. For each scan and method, mean myocardial T1 values were computed for AHA-16 segments. The ICC3 metric was then calculated for all paired test-retest segments to assess the agreement between repeated measurements. A high ICC3 value indicates strong agreement between the test and retest measurements, suggesting robustness and reproducibility in quantifying myocardial T1 values. This metric provides valuable insights into the reliability and consistency of T1 measurements across different methods and scanning conditions. A comprehensive understanding of each method’s performance in clinical practice can be achieved by quantifying the agreement between test and retest measurements across different ROIs and scanning conditions. This analysis also enables the evaluation of the effectiveness of motion correction techniques in improving measurement accuracy and reproducibility.
@@ 5 Results
Figure 3 presents the distribution of the selected reference time points across different acquisition methods (BH-MOLLI, FB-MOLLI, and FB-STONE) by our MBSS-T1 approach. The majority of selections occur at time point 0 across all methods (BH-MOLLI, FB-MOLLI, and FB-STONE). These findings suggest that the reference state is predominantly chosen in the mid-diastolic phase.
@@
Figure 4 depicts the deformed images alongside the predicted deformation fields for the different methods. Our MBSS-T1 approach generates smoother grids with visually realistic deformations, whereas SynthMorph produces less realistic deformation patterns.
Table 2 summarizes the results of the Jacobian determinant analysis. Our MBSS-T1 approach produces deformation fields with a low mean number of folds () and a mean Jacobian determinant close to 1 (), indicating more realistic deformation fields.
| Method | Mean | |